# A perturbative construction of primitive forms from log Landau-Ginzburg mirrors of toric manifolds

### v1

@@ 1. Introduction
The long history of mirror symmetry for toric manifolds can be traced back to the works of Batyrev [4], Givental [19], Lian-Liu-Yau [36], Kontsevich [31], Hori-Vafa [26], and many others. Given an $n$-dimensional projective toric manifold $Y$ defined by a complete smooth fan $\Sigma\subset\mathbb{R}^{n}$, its mirror is generally agreed to be given by a Landau-Ginzburg (abbrev. LG) model $(X_{1},f)$, where $X_{1}$ is a smooth affine variety isomorphic to the algebraic torus (*$(\mathbb{C}^{*})^{n}$ and the so-called superpotential $f:X_{1}\to\mathbb{C}$ is a Laurent polynomial whose Newton polytope is the fan polytope of $\Sigma$. As an example, the LG mirror of $Y=\mathbb{P}^{2}$ is given by $X_{1}=\{(z_{1},z_{2},z_{3})\in\mathbb{C}^{3}:z_{1}z_{2}z_{3}=1\}\cong(\mathbb{C}^{*})^{2}$ together with the restriction $f$ of the function $z_{1}+z_{2}+z_{3}$ to $X_{1}\subset\mathbb{C}^{3}$.
@@
The construction of primitive forms and Frobenius manifold structures is along the lines of the perturbative approach pioneered by Li-Li-Saito [32], which is a complex differential-geometric framework inspired by the BCOV gauge theory of polyvector fields on Calabi-Yau manifolds [5]. Indeed, our results can be viewed as the second example of the general construction in [32], after the example of exceptional unimodular singularities in [33]. More precisely, our primitive forms and Frobenius manifold structures are all determined perturbatively from the central degenerate fiber $X=X_{0}$ of the mirror LG family (or the log LG mirror). This is in sharp contrast with previous constructions, which rely on either the whole or the punctured mirror LG family.
As $X=X_{0}$ is non-compact, we start the construction by following [32] to transform the complex $\operatorname{PV}_{\log}(X)$ of log polyvector fields to that $\operatorname{PV}_{\log,c}(X)$ with compact support via a homotopy (Section 2.3). This allows us to define the higher residue pairing $K_{f}$ on the cohomology $\mathcal{H}_{f}:=H(\operatorname{PV}_{\log}(X)[[u]],\bar{\partial}+df\wedge+u\partial)$ using integration (Section 2.4). Now the main advantage and novelty in choosing the central degenerate fiber $X=X_{0}$ as our reference fiber is that we can choose particularly simple and explicit good bases $\{\varphi_{i}\}$ (Theorem 2.21). In general, the existence of good basis is a highly nontrivial problem. This greatly simplifies the whole construction of the logarithmic Frobenius structure and makes it much more explicit.
@@
At this point, the remaining ingredients in constructing a log Frobenius manifold are a good opposite filtration and a primitive form. The existence of the former is proved using the Deligne extension of a good basis $\{\varphi_{i}\}$ (Lemma 3.8 and Proposition 3.10), while the latter is constructed using the usual Birkhoff factorization procedure (Proposition 3.12). From these we obtain a good opposite filtration and a primitive form $\zeta$, hence yielding a logarithmic Frobenius manifold structure on the base of the universal unfolding and completing the proof of Theorem 1.1; see Section 3.3 for more details.
@@
As our construction is within the framework of Li-Li-Saito [32], the primitive forms and flat coordinates (or semi-infinite period maps) can be computed perturbatively and explicitly. If we write the primitive form as $\zeta=e^{\tilde{F}/u}$, then the corrected superpotential $f+\tilde{F}$, expressed in terms of the flat coordinates, should be closely related to the bulk-deformed Lagrangian Floer superpotential [15, 16] of the projective toric manifold $Y$.
There are very few explicitly known calculations of the bulk-deformed Lagrangian Floer superpotential of a toric manifold. Gross [21, 22] constructed an explicit perturbation of the Hori-Vafa potential $f=z_{1}+z_{2}+z_{3}$ mirror to $Y=\mathbb{P}^{2}$ using counts of Maslov index 2 tropical disks. Recently, Hong-Lin-Zhao [25] provided an inductive algorithm to compute the bulk-deformed Lagrangian Floer superpotentials for Fano toric surfaces using wall-crossing techniques and a tropical-holomorphic correspondence.
@@ Acknowledgement
We would like to thank Hiroshi Iritani, Changzheng Li and Si Li for very helpful discussions.
@@
In this subsection, we explain some notation that will be frequently used. When $P$ is a monoid, $\mathbb{C}[P]$ denotes the monoid ring of $P$ with coefficients in $\mathbb{C}$. We will work with analytic space and write $\text{Spec}(\mathbb{C}[P])$ to be the toric analytic variety. When $X$ is a log analytic space, we will write $X^{\dagger}$ to emphasize the log structure. When $\phi:X^{\dagger}\to Y^{\dagger}$ is a map of log space, $\Omega^{1}_{X^{\dagger}/Y^{\dagger}}$ and $\Theta^{1}_{X^{\dagger}/Y^{\dagger}}$ denote the sheaf of relative log differential and relative log derivation on $X$ respectively. Let $u$ be a formal variable and $A$ be a $\mathbb{C}$-vector space, $A[[u]]$ and $A((u))$ denote the space of $A$-valued formal power series and formal Laurent power series in $u$ respectively.
@@ 2.1. Constructions
In this paper, we assume the reader has some familiarity with log geometry; for references, see [29, 30]. The following is the local model used in the mirror symmetry program of Gross and Siebert, see [24].
Let $N\cong\mathbb{Z}^{n}$ be a lattice of rank $n$ and $N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$ the real vector space it spans. Let $\Sigma\subset N_{\mathbb{R}}$ be a complete smooth fan which defines an $n$-dimensional smooth projective toric manifold $Y$. Since $Y$ is projective, there is a strictly convex piecewise linear function $\psi$ on $|\Sigma|$ with only integral slopes. The subset
@@
The natural inclusion $\mathbb{N}\hookrightarrow P\cap(N\oplus\mathbb{Z})$ given by $q\to(0,q)$ is an inclusion of monoids, inducing a map $\mathscr{X}\to\mathbb{C}=\text{Spec}(\mathbb{C}[\mathbb{N}])$. Let $X$ be the fiber over $0\in\mathbb{C}$. Equip $\mathscr{X}$ with the log structure induced by the inclusion $X\hookrightarrow\mathscr{X}$ and equip $\mathbb{C}$ with the log structure induced by the inclusion $0\hookrightarrow\mathbb{C}$. Then $\Phi:\mathscr{X}^{\dagger}\to\mathbb{C}^{\dagger}$ is log smooth. By pulling back the log structures on $\mathscr{X}^{\dagger}$ and $\mathbb{C}^{\dagger}$ respectively, we get a log smooth map
@@
When $\mathcal{P}$ runs over all primitive collections in $\Sigma(1)$, the equations $\{E_{\mathcal{P}}\}_{\mathcal{P}}$ defines an affine space in $\mathbb{C}^{d}$, and this is exactly the central fiber $X$. In other words, the ideal in $\mathbb{C}[z]:=\mathbb{C}[z_{1},\cdots,z_{d}]$ defining $X$ is the Stenley-Reisner ideal $SR(\Sigma)$ of $\Sigma$. Now $X=\cup_{i}X_{i}$ is the union of $n$-dimensional coordinate hyperplanes in $\mathbb{C}^{d}$, each corresponding to an $n$-dimensional cone of $\Sigma$. For convenience, we denote by $X_{1}$ the component defined by $z_{l}=0$ for $n+1\leq l\leq d$.
Since $\phi:X^{{\dagger}}\to 0^{\dagger}$ is log smooth, the sheaf of (relative) log differentials $\Omega^{\bullet}_{X^{\dagger}/0^{\dagger}}$ and the sheaf of (relative) log derivations $\Theta^{\bullet}_{X^{\dagger}/0^{\dagger}}$ are both locally free on $X$. Let us describe these sheaves explicitly. Since $\Sigma$ is smooth, we can assume, without loss of generality, that $\rho_{1}=e_{1},\cdots,\rho_{n}=e_{n}$ is the standard basis of $N$. Assume that when $n+1\leq l\leq d$,
@@
We set $\Omega^{k}_{X^{\dagger}/0^{\dagger}}:=\wedge^{k}\Omega^{1}_{X^{\dagger}/0^{\dagger}}$. On any $n$-dimensional component $X_{i}$ of $X$, any $\alpha\in\Omega^{k}_{X^{\dagger}/0^{\dagger}}$ has a unique representative $\alpha|_{X_{i}}$ given by a holomorphic $k$-form with at most log poles on $X_{j}\cap\text{Sing}\,X$. The compatibility condition is that if $X_{i}$ and $X_{j}$ has $(n-1)$-dimensional intersection $D_{ij}$, then the restriction of $\alpha|_{X_{i}}$ and $\alpha|_{X_{j}}$ to $D_{ij}$ agrees when taking into account the relations (1).
@@
[…]
@@ 2.2. The state space isomorphism
Note that $\{f,-\}$ acts as a derivation, hence the cohomology group has a product structure. Let $u$ be a formal variable. Then we can consider
@@
where $[[u]]$ denotes formal power series in $u$. Again by conjugating with $\Omega$, we obtain
@@
Using the conjugation by $\Omega$, the theorem then follows from the well-known result on the cohomologies of compact toric manifolds. In the equation (3), $P(\Sigma)$ denotes the ideal in $\mathbb{C}[z]$ generated by $n$ elements
@@
Since $H^{*}(\Theta^{\bullet}_{X^{\dagger}/0^{\dagger}},\{f,-\})$ is concentrated at degree 00, we have the following Hodge-to-de Rham degeneration property:
@@ 2.3. A homotopy formula
Let $\mathcal{A}^{0,j}$ be the sheaf of smooth differential $(0,j)$-forms on $X$ and define
@@
00 0 $\to\Omega^{p}_{X^{\dagger}/0^{\dagger}}\to\mathcal{A}^{p,0}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\mathcal{A}^{p,1}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\mathcal{A}^{p,2}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\cdots,$
00 0 $\to\Theta^{p}_{X^{\dagger}/0^{\dagger}}\to\operatorname{PV}^{p,0}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\operatorname{PV}^{p,1}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\operatorname{PV}^{p,2}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\cdots.$
@@
and denote by $\mathcal{A}_{\log,c}(X)$ and $\operatorname{PV}_{\log,c}(X)$ the subspaces of the corresponding spaces that consist of elements with only compact support. It is clear that they are all super-commutative algebras over the algebra of smooth functions on $X$. In the sequel, if $\alpha$ is in $\mathcal{A}^{i,j}_{\log}(X)$ or $\operatorname{PV}^{i,j}_{\log}(X)$, we will call it a differential form or polyvector field of bidgree $(i,j)$, and define $|\alpha|=i+j$ for $\alpha\in\mathcal{A}^{i,j}_{\log}(X)$ and $|\alpha|=i-j$ for $\alpha\in\operatorname{PV}^{i,j}_{\log}(X)$. By abuse of notation, we will use ¯$\bar{\partial}_{f}$ to denote the operator ¯$\bar{\partial}+\text{d}f\wedge$ when it acts on differential forms and the operator ¯+$\bar{\partial}+\{f,-\}$ when it acts on polyvector fields.
@@
$H(\Omega^{\bullet}_{X^{\dagger}/0^{\dagger}},\{f,-\})\cong$ $H(\mathcal{A}_{\log}(X),\bar{\partial}_{f}),$
@@
If we denote by $Q_{f}$ the operator ¯$\bar{\partial}_{f}+u\partial$, we have a similar result:
@@
[…]
@@
Let $\rho=\rho(|z_{1}|^{2},\cdots,|z_{d}|^{2})$ be a smooth (real-valued) cut-off function with compact support such that $\rho\equiv 1$ in a neighbourhood of 00. For $1\leq k\leq n$, write $g_{k}:=\sum_{i=1}^{d}\langle\rho_{i},e^{k}\rangle z_{i}=z_{k}+\sum_{l=n+1}^{d}a_{kl}z_{l}$, and define
@@
of $\Theta_{X^{\dagger}/0^{\dagger}}^{1}$, we have $V_{f}=\frac{1}{\sum_{k=1}g_{k}\bar{g}_{k}}\left(\sum_{k=1}^{n}\bar{g}_{k}\tilde{\theta}_{k}\right)$. The zeros of $\sum_{k=1}g_{k}\bar{g}_{k}$ contain only the origin since $\Sigma$ is smooth. As $V_{f}$ satisfies the condition of (2), $V_{f}$ is a relative log derivation on $X-0$ and it defines an operator acting on $\operatorname{PV}_{\log}(X-0)$.
@@
By conjugating with $\Omega$ and using the frame $(\tilde{\theta}_{i_{1}},\cdots,\tilde{\theta}_{i_{n}})=(\tilde{\theta}_{1},\cdots,\tilde{\theta}_{n})(C^{T})^{-1}$,
$\{f,-\}=\left(\frac{\partial}{\partial\tilde{\theta}_{i_{1}}},\cdots,\frac{\partial}{\partial\tilde{\theta}_{i_{n}}}\right)(C_{i})^{-1}(g_{1},\cdots,g_{n})^{T},$
where $[\frac{\partial}{\partial\tilde{\theta}_{i_{j}}},\tilde{\theta}_{i_{s}}\wedge]=\delta_{js}$ as operators. On the other hand, on $X_{i}-0$ we can write
@@
Hence [$[\{f,-\},V_{f}]=\frac{1}{\sum_{k=1}g_{k}\bar{g}_{k}}(\bar{g}_{1},\cdots,\bar{g}_{k})C_{i}(C_{i})^{-1}(g_{1},\cdots,g_{n})^{T}=1$. ∎

### v3

@@ 1. Introduction
The long history of mirror symmetry for toric manifolds can be traced back to the works of Batyrev [4], Givental [19], Lian-Liu-Yau [36], Kontsevich [31], Hori-Vafa [26], and many others. Given an $n$-dimensional projective toric manifold $Y$ defined by a complete smooth fan $\Sigma\subset\mathbb{R}^{n}$, its mirror is generally agreed to be given by a Landau-Ginzburg (abbrev. LG) model $(X_{1},f)$, where $X_{1}$ is a smooth affine variety isomorphic to the algebraic torus $(\mathbb{C}^{*})^{n}$ and the so-called superpotential $f:X_{1}\to\mathbb{C}$ is a Laurent polynomial whose Newton polytope is the fan polytope of $\Sigma$. As an example, the LG mirror of $Y=\mathbb{P}^{2}$ is given by $X_{1}=\{(z_{1},z_{2},z_{3})\in\mathbb{C}^{3}:z_{1}z_{2}z_{3}=1\}\cong(\mathbb{C}^{*})^{2}$ together with the restriction $f$ of the function $z_{1}+z_{2}+z_{3}$ to $X_{1}\subset\mathbb{C}^{3}$.
@@
The construction of primitive forms and Frobenius manifold structures is along the lines of the perturbative, explicit approach using polyvector fields pioneered by Li-Li-Saito [32] (see also [41, 42]), which is a complex differential-geometric framework inspired by the BCOV gauge theory on Calabi-Yau manifolds [5]. Indeed, our results can be viewed as the second application of the general construction and explicit perturbative formula in [32], after that of exceptional unimodular singularities in [33]. More precisely, our primitive forms and Frobenius manifold structures can all be determined perturbatively and explicitly from the central degenerate fiber $X=X_{0}$ of the mirror LG family (or the log LG mirror). This is in sharp contrast with previous constructions, which rely on either the entire or punctured mirror LG family.
As $X=X_{0}$ is non-compact, we start the construction by following [32] to transform the complex $\operatorname{PV}_{\log}(X)$ of log polyvector fields to that $\operatorname{PV}_{\log,c}(X)$ with compact support via a homotopy (Section 2.3). This allows us to define the higher residue pairing $K_{f}$ on the (formally completed) Brieskorn lattice $\mathcal{H}_{f}:=H(\operatorname{PV}_{\log}(X)[[u]],\bar{\partial}+df\wedge+u\partial)$ using integration (Section 2.4). Now the main advantage and novelty in choosing the central degenerate fiber $X=X_{0}$ as our reference fiber is that we can find particularly simple and explicit good bases $\{\varphi_{i}\}$ (Theorem 2.21) via the state space isomorphism (Theorem 2.4). In general, the existence of good basis is a highly nontrivial problem. This greatly simplifies the construction of the logarithmic Frobenius structure.
@@
At this point, the remaining ingredients in constructing a log Frobenius manifold are a good opposite filtration and a primitive form. The existence of the former is proved using the Deligne extension of a good basis $\{\varphi_{i}\}$ (Lemma 3.8 and Proposition 3.10), while the latter is constructed using the standard Birkhoff factorization procedure (Equation (15), Lemma 3.11, Proposition 3.12). From these we obtain a good opposite filtration and a primitive form $\zeta$, hence yielding a logarithmic Frobenius manifold structure on the base of the universal unfolding and completing the proof of Theorem 1.1; see Section 3.3 for more details.
@@
As our construction is within the framework of Li-Li-Saito [32], the primitive forms and flat coordinates (or semi-infinite period maps) can all be computed perturbatively and explicitly. If we write the primitive form as $\zeta=e^{\tilde{F}/u}$, then the corrected superpotential $f+\tilde{F}$, expressed in terms of the flat coordinates (and with the descendant variable set to be zero), should be closely related to the bulk-deformed Lagrangian Floer superpotential [15, 16] of the projective toric manifold $Y$.
There are very few explicitly known calculations of the bulk-deformed Lagrangian Floer superpotential of a toric manifold. Gross [21, 22] constructed an explicit perturbation of the Hori-Vafa potential $f=z_{1}+z_{2}+z_{3}$ mirror to $Y=\mathbb{P}^{2}$ using counts of Maslov index 2 tropical disks. Recently, Hong-Lin-Zhao [25] generalized this and provided an inductive algorithm to compute the bulk-deformed Lagrangian Floer superpotentials for all Fano toric surfaces using wall-crossing techniques and a tropical-holomorphic correspondence.
@@
Remark 1.2.
The explicit computations in Section 3.4.2 are all done by hand. In general, the semi-infinite period map and primitive form (and hence the corrected superpotential) of any log LG toric mirror can all be computed up to any desired order using a computer program.
@@ Acknowledgement
We would like to thank Hiroshi Iritani, Changzheng Li, Si Li and Junwu Tu for very helpful discussions. K. Chan was supported by grants from the Hong Kong Research Grants Council (Project No. CUHK14301420, CUHK14301621 & CUHK14305322). Z. N. Ma was supported by National Natural Science Foundation of China (Grant No. K23281001, K23281103). H. Wen was supported by the Young Scientists Fund of the National Natural Science Foundation of China (Grant No. 12201314).
@@
In this subsection, we explain some notation that will be frequently used. When $P$ is a monoid, $\mathbb{C}[P]$ denotes the monoid ring of $P$ with coefficients in $\mathbb{C}$. We will work with analytic spaces and write $\text{Spec}(\mathbb{C}[P])$ as the corresponding toric analytic variety. When $X$ is a log analytic space, we will write $X^{\dagger}$ to emphasize its log structure. When $\phi:X^{\dagger}\to Y^{\dagger}$ is a map of log spaces, $\Omega^{1}_{X^{\dagger}/Y^{\dagger}}$ and $\Theta^{1}_{X^{\dagger}/Y^{\dagger}}$ denote the sheaf of relative log differentials and sheaf of relative log derivations on $X$ respectively. Let $u$ be a formal variable and $A$ be a $\mathbb{C}$-vector space. Then $A[[u]]$ and $A((u))$ denote the space of $A$-valued formal power series and formal Laurent power series in $u$ respectively.
@@ 2.1. Constructions
In this paper, we assume the reader has some familiarity with log geometry; for references, see [29, 30]. We first describe the local model used in the mirror symmetry program of Gross and Siebert, see [24].
Let $N\cong\mathbb{Z}^{n}$ be a lattice of rank $n$ and $N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$ the real vector space it spans. Let $\Sigma\subset N_{\mathbb{R}}$ be a smooth complete fan which defines an $n$-dimensional smooth projective toric manifold $Y$. Since $Y$ is projective, there is a strictly convex piecewise linear function $\psi$ on $|\Sigma|$ with only integral slopes. The subset
@@
The natural inclusion $\mathbb{N}\hookrightarrow P\cap(N\oplus\mathbb{Z})$ given by $q\to(0,q)$ is an inclusion of monoids, which induces a map $\mathscr{X}\to\mathbb{C}=\text{Spec}(\mathbb{C}[\mathbb{N}])$. Let $X$ be the fiber over $0\in\mathbb{C}$. Equip $\mathscr{X}$ with the log structure induced by the inclusion $X\hookrightarrow\mathscr{X}$ and equip $\mathbb{C}$ with the log structure induced by the inclusion $0\hookrightarrow\mathbb{C}$. Then $\Phi:\mathscr{X}^{\dagger}\to\mathbb{C}^{\dagger}$ is log smooth. By pulling back the log structures on $\mathscr{X}^{\dagger}$ and $\mathbb{C}^{\dagger}$ respectively, we get a log smooth map
@@
When $\mathcal{P}$ runs over all primitive collections in $\Sigma(1)$, the equations $\{E_{\mathcal{P}}\}_{\mathcal{P}}$ define an affine space in $\mathbb{C}^{d}$, and this is exactly the central fiber $X$. In other words, the ideal in $\mathbb{C}[z]:=\mathbb{C}[z_{1},\cdots,z_{d}]$ defining $X$ is the Stenley-Reisner ideal $SR(\Sigma)$ of $\Sigma$. Now $X=\cup_{i}X_{i}$ is the union of $n$-dimensional coordinate hyperplanes in $\mathbb{C}^{d}$, each corresponding to an $n$-dimensional cone of $\Sigma$. For convenience, we denote by $X_{1}$ the component defined by $z_{l}=0$ for $n+1\leq l\leq d$.
Since $\phi:X^{{\dagger}}\to 0^{\dagger}$ is log smooth, the sheaf $\Omega^{\bullet}_{X^{\dagger}/0^{\dagger}}$ of (relative) log differentials and the sheaf $\Theta^{\bullet}_{X^{\dagger}/0^{\dagger}}$ of (relative) log derivations are both locally free on $X$. Let us describe these sheaves explicitly. Since $\Sigma$ is smooth, we can assume, without loss of generality, that $\rho_{1}=e_{1},\cdots,\rho_{n}=e_{n}$ is the standard basis of $N$. Assume that when $n+1\leq l\leq d$,
@@
We set $\Omega^{k}_{X^{\dagger}/0^{\dagger}}:=\wedge^{k}\Omega^{1}_{X^{\dagger}/0^{\dagger}}$. On any $n$-dimensional component $X_{i}$ of $X$, any $\alpha\in\Omega^{k}_{X^{\dagger}/0^{\dagger}}$ has a unique representative $\alpha|_{X_{i}}$ given by a holomorphic $k$-form with at most log poles on $X_{j}\cap\text{Sing}\,X$. The compatibility condition is that if $X_{i}$ and $X_{j}$ has an (n1)(n-1)-dimensional intersection $D_{ij}$, then the restriction of $\alpha|_{X_{i}}$ and $\alpha|_{X_{j}}$ to $D_{ij}$ agrees when taking into account the relations (1).
@@
A logarithmic Landau-Ginzburg model (abbrev. log LG model) is a triple $(X^{\dagger},\phi,f)$, where $X^{\dagger}$ is a log Calabi-Yau analytic space with underlying space $X$, $\phi:X^{\dagger}\to 0^{\dagger}$ is a log smooth map from $X^{\dagger}$ to the standard $P$-log point $0^{\dagger}$ associated to some toric monoid $P$, and $f$ is a holomorphic function (called the superpotential) on $X$ such that when $\text{d}f$ is viewed as a relative log differential, the critical set
$\text{Crit}(f):=\{z\in\mathbb{C}^{d}:\text{d}f(z)=0\}$
@@ 2.2. The state space isomorphism
Note that $\{f,-\}$ acts as a derivation, hence the cohomology group has a product structure.
Let $u$ be a formal variable. Then we can consider the complex
@@
where $[[u]]$ denotes formal power series in $u$. Again by conjugating with $\Omega$, we obtain another complex
@@
Using the conjugation by $\Omega$, the theorem then follows from the well-known result on the cohomologies of compact toric manifolds. In equation (3), $P(\Sigma)$ denotes the ideal in $\mathbb{C}[z]$ generated by $n$ elements
@@
Since $H^{*}(\Theta^{\bullet}_{X^{\dagger}/0^{\dagger}},\{f,-\})$ is concentrated at degree 0, we have the following Hodge-to-de Rham degeneration property:
@@ 2.3. A homotopy formula
Let $\mathcal{A}^{0,j}$ be the sheaf of smooth differential $(0,j)$-forms on $X$. Concretely, $\alpha\in\mathcal{A}^{0,j}(U)$ is given by a smooth $(0,j)$-form $\alpha_{i}$ on each $U\cap X_{i}$ such that the pull-back of $\alpha_{i}$ and $\alpha_{j}$ to $U\cap X_{i}\cap X_{j}$ coincide. Define
@@
$0$ $\to\Omega^{p}_{X^{\dagger}/0^{\dagger}}\to\mathcal{A}^{p,0}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\mathcal{A}^{p,1}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\mathcal{A}^{p,2}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\cdots,$
$0$ $\to\Theta^{p}_{X^{\dagger}/0^{\dagger}}\to\operatorname{PV}^{p,0}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\operatorname{PV}^{p,1}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\operatorname{PV}^{p,2}_{\log}\stackrel{{\scriptstyle\bar{\partial}}}{{\longrightarrow}}\cdots.$
@@
and denote by $\mathcal{A}_{\log,c}(X)$ and $\operatorname{PV}_{\log,c}(X)$ the subspaces of the corresponding spaces that consist of elements with only compact support. It is clear that they are all super-commutative algebras over the algebra of smooth functions on $X$. In the sequel, if $\alpha$ is in $\mathcal{A}^{i,j}_{\log}(X)$ or $\operatorname{PV}^{i,j}_{\log}(X)$, we will call it a differential form or polyvector field of bidgree $(i,j)$, and define $|\alpha|=i+j$ for $\alpha\in\mathcal{A}^{i,j}_{\log}(X)$ and $|\alpha|=i-j$ for $\alpha\in\operatorname{PV}^{i,j}_{\log}(X)$. By abuse of notation, we will use $\bar{\partial}_{f}$ to denote the operator $\bar{\partial}+\text{d}f\wedge$ when it acts on differential forms and the operator $\bar{\partial}+\{f,-\}$ when it acts on polyvector fields.
@@
$H(\Omega^{\bullet}_{X^{\dagger}/0^{\dagger}},\text{d}f)\cong$ $H(\mathcal{A}_{\log}(X),\bar{\partial}_{f}),$
@@
If we denote by $Q_{f}$ the operator $\bar{\partial}_{f}+u\partial$, we have a similar result:
@@
$H(\Omega^{\bullet}_{X^{\dagger}/0^{\dagger}}[[u]],\text{d}f\wedge+u\partial)\cong$ $H(\mathcal{A}_{\log}(X)[[u]],Q_{f}),$
@@
Let $\rho=\rho(|z_{1}|^{2},\cdots,|z_{d}|^{2})$ be a smooth (real-valued) cut-off function with compact support such that $\rho\equiv 1$ in a neighbourhood of 0. For $1\leq k\leq n$, write $g_{k}:=\sum_{i=1}^{d}\langle\rho_{i},e^{k}\rangle z_{i}=z_{k}+\sum_{l=n+1}^{d}a_{kl}z_{l}$, and define
@@
of $\Theta_{X^{\dagger}/0^{\dagger}}^{1}$, we have $V_{f}=\frac{1}{\sum_{k=1}g_{k}\bar{g}_{k}}\left(\sum_{k=1}^{n}\bar{g}_{k}\tilde{\theta}_{k}\right)$. The zeros of $\sum_{k=1}g_{k}\bar{g}_{k}$ contain only the origin since $\Sigma$ is smooth. As $V_{f}$ satisfies the condition of (2), $V_{f}$ is a relative log derivation on X0X-0 and it defines an operator acting on $\operatorname{PV}_{\log}(X-0)$.
@@
By conjugating with $\Omega$ and using the frame $(\tilde{\theta}_{i_{1}},\cdots,\tilde{\theta}_{i_{n}})=(\tilde{\theta}_{1},\cdots,\tilde{\theta}_{n})(C_{i}^{T})^{-1}$,
(6) $\{f,-\}=\left(\frac{\partial}{\partial\tilde{\theta}_{i_{1}}},\cdots,\frac{\partial}{\partial\tilde{\theta}_{i_{n}}}\right)(C_{i})^{-1}(g_{1},\cdots,g_{n})^{T},$
where $\frac{\partial}{\partial\tilde{\theta}_{i_{j}}},\tilde{\theta}_{i_{s}}\wedge]=\delta_{js}$ as operators. On the other hand, on $X_{i}-0$ we can write
@@
Hence $\{f,-\},V_{f}]=\frac{1}{\sum_{k=1}g_{k}\bar{g}_{k}}(\bar{g}_{1},\cdots,\bar{g}_{k})C_{i}(C_{i})^{-1}(g_{1},\cdots,g_{n})^{T}=1$. ∎